OpenStudy (anonymous):

Use the quadratic formula to solve the equation. If necessary, round to the nearest hundredth. x^2+3=-4x

OpenStudy (shayaan_mustafa):

Hi again :) Do you know what is quadratic formula dear ?

OpenStudy (anonymous):

answers: a. 1,3 b. -1, -3 c. 1, -3

OpenStudy (anonymous):

no i dont

OpenStudy (shayaan_mustafa):

\[\huge x=\frac{-b \pm \sqrt{b^2-4ac}}{2a}\] Where your a= coefficient of your x^2 term b= coefficient of your x term c= constant

OpenStudy (anonymous):

so it would be 1, 3?

OpenStudy (shayaan_mustafa):

No.

OpenStudy (shayaan_mustafa):

Now look carefully I tell you one shortcut to get result. Are you present here?

OpenStudy (shayaan_mustafa):

@bodetheenglishmastiff

OpenStudy (anonymous):

yes

OpenStudy (shayaan_mustafa):

It is your equation \[\large x^2+3=-4x\] OK. Now you have 3 choices either of one is correct. Now substitute one by one in the equation and check whether choice is correct or not.If choice is correct then left hand side and right hand side must be equal. (1,3) \[\large (1)^2+3=-4(1)\] \[\large 1+3=-4\] \[\large 4=-4\] As you can see both sides are not equal. So we will not repeat same thing with three in this choice. Now consider (-1,-3) \[\large (-1)^2+3=-4(-1)\] \[\large 1+3=4\] \[\large 4=4\] Now same for -3 \[\large (-3)^2+3=-4(-3)\] \[\large 9+3=12\] \[\large 12=12\] As you can see both sides are equal by substituting -1 and -3. So b.(-1,-3) is your answer Regards Electronics Engineer

OpenStudy (shayaan_mustafa):

Similarly you can check your choice c.(1,-3)

OpenStudy (anonymous):

your amazing! thanks soooo much!!!!

OpenStudy (shayaan_mustafa):

My pleasure. :D We play with maths every day.. LOL